# Slack-Pack algorithm for Meir-Moser packing problem

### v2

@@ Abstract
The well-known problem stated by A. Meir and L. Moser consists in tiling the unit square with rectangles (details), whose side lengths equal $\frac{1}{n}\times\frac{1}{n+1}$, where indices $n$ range from 1 to infinity. Recently, Terence Tao has proved that it is possible to tile with $\bigl{(}\frac{1}{n}\bigr{)}^{t}\times\bigl{(}\frac{1}{n+1}\bigr{)}^{t}$ rectangles (squares with the side length of $\bigl{(}\frac{1}{n}\bigr{)}^{t}$), $11/2<t<$ 1, the square, whose area equals the sum of areas of these details, provided that only those details, whose indices exceed certain $n_{0}$, are taken into consideration. We adduce arguments in favor of the assumption that the result obtained by T. Tao is also valid for $t=1$. We use a new tiling method (the Slack-Pack algorithm), which initially admits gaps between stacks of details. The algorithm uses a pre-fixed parameter $\gamma$, $\sqrt{3/2}<\gamma<3/2$, connected with the gap value. The new algorithm allows one to control the ratio of the area of the large rectangular part, which is free of details, to the whole area of the remaining empty space. This ratio (under certain natural assumptions) always exceeds $1-1/\gamma-\delta$, where $\delta$ tends to zero as $n_{0}$ increases.
@@ 2.2 The Slack-Pack algorithm
This algorithm allows us to perfectly pack both rectangular and square details. The input data for the algorithm is $n_{0}$, i.e., the index of the first detail that has to be packed. Let the symbol $n$ denote the current time moment, and let $\gamma$ stand for the parameter used by the algorithm, $\sqrt{3/2}<\gamma<3/2$.
@@ 2.3 The main theorem
Let the symbol $t_{0}$ stand for the starting moment of the second stage (if it ever starts). Later we will prove (see Lemma 8) that $t_{0}\sim n_{0}^{\gamma}$, where the notation $\sim$ means that the limit value of the ratio of sequences equals one. Evidently, the onset of the time moment $t_{0}$ does not necessarily imply the infinity of the algorithm operation. The necessary and sufficient condition for it is the possibility of the implementation of Step 4, i.e., the sufficiency of sizes of the LRP for cutting the next stripe. In what follows, we also use the symbol $\sim$ for comparing the behavior of other parameters of the algorithm with $n_{0}\to\infty$.
@@
Evidently, this condition ensures the possibility of a perfect packing. We treat such time moments as critical and denote them as $T_{crit}$. Note that at these time moments, when we have to cut a stripe from the LRP, the following inequality is fulfilled:
@@
In the next section, we prove one important property of shapes of normal boxes, namely, For all $B\in{\mathcal{B}}_{norm,1}$, with any arbitrarily small $\sigma>0$ and sufficiently large $n_{0}$, the following inequality is valid:
@@
The following proposition is the key lemma in this paper:
@@
Assume that the width of the endpoint obtained at time moment $n$ represents a random value $\xi_{n}$ uniformly distributed on the segment, whose left and right boundaries have, correspondingly, the order of the minimal and maximal possible width of an endpoint at time moment $n$, while the order of the height of this endpoint is $1/n$. Note that at time moment $t$ we have a set of endpoints obtained at time moments $n$, $n<t$. Assume that they obey a probability distribution law which agrees with all previously made assumptions. In Appendix A, we prove that for such a mixture of uniformly distributed random values $\xi_{n}$, the asymptotics of the mean value of the expression $(1/n)/\xi_{n}$ is $(\gamma^{2}-1)\ln(t)$. We also adduce results of numerical experiments which with various $\gamma$ prove that the following ratio of the sample average to $\ln(t)$
$\frac{1}{\ln(t)\,|{\mathcal{B}}_{ep,2}(t)|}\sum_{B\in{\mathcal{B}}_{ep,2}(t)}\frac{h(B)}{w(B)}.$
@@ 3 Lemmas on normal boxes
Formula (3) implies that for any $\delta>0$ with sufficiently large $n_{0}$ the following inequality is valid for all $B\in{\mathcal{B}}(t)$, $t\in T_{crit}$:
@@ 4.2 Lemmas on the starting time moment of the second stage of the algorithm
Therefore, $k_{0}\sim\sqrt{n_{0}}$ and the difference of widths $\Delta$ of the first detail in the stripe and the last one is $\Delta\sim 1/n_{0}-1/(n_{0}+k_{0})\sim n_{0}^{-3/2}$. Let the symbol $t^{\prime}$ stand for the first time moment, when the next detail can be packed in the box ${\widetilde{B}}$ with the necessary gap, whose order of magnitude is $(t^{\prime})^{-\gamma}$. Then $1/t^{\prime}+(t^{\prime})^{-\gamma}\geq 1/n_{0}+\Delta$, and since $\gamma^{2}>3/2$, we conclude that $t^{\prime}\sim n_{0}^{\gamma}$.
@@
The starting time moment of the second stage $t_{0}$ can be somewhat larger than $t^{\prime}$ for two reasons: either due to the delay of the order of magnitude of $O(\sqrt{t^{\prime}})$ caused by the fact that the cut-off stripe is already full with details, or due to the existence of a large number of endpoints, whose width exceeds $w({\widetilde{B}})$. Since the order of magnitude of the value added due to the first reason is only $O(\sqrt{t^{\prime}})$, which is asymptotically neglectable in comparison with $t^{\prime}$, it remains to consider only the second reason.
Assume that we can pack in endpoints only details with indices ranging from $t^{\prime}$ to $Ct^{\prime}$, where $C>1$ is some fixed number. But then the total area of these boxes is asymptotically not less than
@@
Really, inequality (4) is equivalent to the fact that the box $D_{n}$ satisfies the inequality $(1+\sigma)/n^{\gamma}>w(D_{n})$, therefore the width of any normal box exceeds $(1+\sigma)/n_{0}^{\gamma}$, whence we conclude that $t_{0}>n_{0}^{\gamma}/(1+\sigma)$. Two-sided constraints imply the correlation $t_{0}\sim n_{0}^{\gamma}$.
@@ 5 Proof of the main theorem
$S_{ep,2}(t)=\sum_{B\in{\mathcal{B}}_{ep,2}}w^{2}(B)\frac{h(B)}{w(B)}=O\bigl{(}\frac{1}{t^{2}}\bigr{)}|{\mathcal{B}}_{ep,2}|\times\frac{1}{|{\mathcal{B}}_{ep,2}(t)|}\sum_{B\in{\mathcal{B}}_{ep,2}(t)}\frac{h(B)}{w(B)}.$
@@
In addition, according to formula (7), $S_{com}(t)\sim 1/t$. For $t\leq t_{0}$ the assertion of the theorem is valid by Lemma 8, and there is no need in using assumptions 1 and 2. Let us now consider the general case for all $t\in T_{crit}$. Let us divide the left- and right-hand sides of equality (13) by $S_{com}(t)$. The second addend satisfies inequality (5), the third one does Lemma 4, the fourth one, according to Lemma 10, satisfies the bound $S_{ep,2}(t)/S_{com}(t)=o(t)$. ∎
@@ 6 Conclusion
The graphs given below illustrate the dependence of the ratio of LRP on the value of the parameter $n_{0}$ with $\gamma=4/3$, when the Slack-Pack algorithm is being used for packing up to $10^{10}$ details. The horizontal dashed line corresponds to the value of $1-1/\gamma$, while the vertical dash-dotted line does to the argument $n_{0}^{\gamma}\sim t_{0}$. According to experimental results, after attaining the value $t=t_{0}$ the ratio of LRP stabilizes at the level slightly exceeding $1-1/\gamma-\delta$, where $\delta$ is some small value.
@@
We have also performed certain experiments for verifying inequality (4) in Assumption 1. For the value $1+\sigma$ we used the expression $h({\widetilde{B}})^{\gamma}/w({\widetilde{B}})$, where ${\widetilde{B}}$ denotes the last normal box in the first stripe. Experiments were performed for sets of details $R_{n}$ with the same values of $n_{0}$ and $\gamma$ as those used in previous experiments. According to obtained results, inequality (4) is fulfilled for all considered normal boxes. See also the Appendix A for theoretical and experimental arguments in favor of Assumption 2.

### v3

@@ Abstract
The well-known problem stated by A. Meir and L. Moser consists in tiling the unit square with rectangles (details), whose side lengths equal $1/n\times 1/(n+1)$, where indices $n$ range from 1 to infinity. Recently, Terence Tao has proved that it is possible to tile with $1/n^{t}\times 1/(n+1)^{t}$ rectangles (squares with the side length of $1/n^{t}$), $11/2<t<$ 1, the square, whose area equals the sum of areas of these details, provided that only those details, whose indices exceed certain $n_{0}$, are taken into consideration. We adduce arguments in favor of the assumption that the result obtained by T. Tao is also valid for $t=1$. We use a new tiling method (the Slack-Pack algorithm), which initially admits gaps between stacks of details. The algorithm uses a pre-fixed parameter $\gamma$, $\sqrt{3/2}<\gamma<3/2$, connected with the gap value. The new algorithm allows one to control the ratio of the area of the large rectangular part, which is free of details, to the whole area of the remaining empty space. This ratio (under certain natural assumptions) always exceeds $1-1/\gamma-\delta$, where $\delta$ tends to zero as $n_{0}$ increases.
@@ 2.2 The Slack-Pack algorithm
This algorithm allows us to perfectly pack both rectangular and square details. The input data for the algorithm is $n_{0}$, i.e., the index of the first detail that has to be packed. Let the symbol $n$ denote the current time moment, and let $\gamma$ stand for the parameter used by the algorithm, $\sqrt{3/2}<\gamma<3/2$. We will also use the parameter $N$ to control the size of gaps between packed details.
@@ 2.3 The main theorem
Let the symbol $t_{0}$ stand for the starting moment of the second stage (if it ever starts). Later we will prove (see Lemma 8) that $t_{0}\sim n_{0}^{\gamma}$, where the notation $\sim$ means that the limit of the ratio $t_{0}/n_{0}^{\gamma}$ viewed as a function of $n_{0}$ tends to $1$ as $n_{0}\to\infty$. Evidently, the onset of the time moment $t_{0}$ does not necessarily imply the infinity of the algorithm operation. The necessary and sufficient condition for it is the possibility of the implementation of Step 4, i.e., the sufficiency of sizes of the LRP for cutting the next stripe. In what follows, we also use the symbol $\sim$ for comparing the behavior of other parameters of the algorithm with $n_{0}\to\infty$.
@@
We treat such time moments as critical and denote them as $T_{crit}$.
Inequality (2) states that $w(LRP(t))\times h(LRP(t))/S_{com}(t)>c>0$ for all moments $t\in T_{crit}$. It is straightforward to show that in this case $w(LRP(t))=\Theta(1/\sqrt{t})$ and $h(LRP(t))=\Theta(1/\sqrt{t})$ if $n_{0}$ is chosen sufficiently large. Hence, the detail $D_{t}$ will successfully fit in Step 4 into the LRP since sizes of the detail are $o(1/\sqrt{t})$ and algorithm will not stop.
This consequence of the inequality $S_{LRP}(t)/S_{com}(t)>c>0$ is also used in other algorithms where LRP arises, therefore the study of the asymptotic behavior of the fraction $S_{LRP}(t)/S_{com}(t)$ is extremely important.
Note that at the time moments, when we have to cut a stripe from the LRP, the following inequality is fulfilled:
@@
For details $R_{n}$ inequality (3) can be strengthened by replacing $1/t$ with $1/(t+1)$. However, for our subsequent analysis, it is more convenient to keep it in this universal form.
@@
In the next section, we prove one important property of shapes of normal boxes, namely, For all $B\in{\mathcal{B}}_{norm,1}$, with any arbitrarily small $\sigma>0$ and sufficiently large $n_{0}$ (where $n_{0}$ depends on $\sigma$), the following inequality is valid:
@@
The following lemma is the key lemma in this paper:
@@
Assume that the width of the endpoint obtained at time moment $n$ represents a random value $\xi_{n}$ uniformly distributed on the segment, whose left and right boundaries have, correspondingly, the order of the minimal and maximal possible width of an endpoint at time moment $n$, while the order of the height of this endpoint is $1/n$. Note that at time moment $t$ we have a set of endpoints obtained at time moments $n$, $n<t$. Assume that they obey a probability distribution law which agrees with all previously made assumptions. In Appendix A, we prove that for such a mixture of uniformly distributed random values $\xi_{n}$, the asymptotics of the mean value of the expression $(1/n)/\xi_{n}$ is $(\gamma^{2}-1)\ln(t)$. We also adduce results of numerical experiments which, for various values of $\gamma$, allow us to assume that the following ratio of the sample average to $\ln(t)$
$\frac{1}{\ln(t)\,|{\mathcal{B}}_{ep,2}(t)|}\sum_{B\in{\mathcal{B}}_{ep,2}(t)}\frac{h(B)}{w(B)}$
@@ 3 Lemmas on normal boxes
Formula (3) implies that for any $\delta>0$ with $n_{0}>({1}/{\delta})^{\gamma-1}$ the following inequality is valid for all $B\in{\mathcal{B}}(t)$, $t\in T_{crit}$:
@@ 4.2 Lemmas on the starting time moment of the second stage of the algorithm
Therefore, $k_{0}\sim\sqrt{n_{0}}$ and the difference of widths $\Delta$ of the first detail in the stripe and the last one is $\Delta\sim 1/n_{0}-1/(n_{0}+k_{0})\sim n_{0}^{-3/2}$. Let the symbol $t^{\prime}$ stand for the first time moment, when the next detail can be packed in the box ${\widetilde{B}}$ with the necessary gap, whose order of magnitude is $(t^{\prime})^{-\gamma}$. Then $1/t^{\prime}+(t^{\prime})^{-\gamma}\geq(1/n_{0})^{\gamma}+\Delta$, and since $\gamma^{2}>3/2$, we conclude that $t^{\prime}\sim n_{0}^{\gamma}$.
@@
The starting time moment of the second stage $t_{0}$ can be somewhat larger than $t^{\prime}$ for two reasons. First, there might be a delay of order $O(\sqrt{t^{\prime}})$ caused by the cut-off stripe still being filled with details. The second possible reason is the existence of a large number of endpoints, whose width exceeds $w({\widetilde{B}})$. Since the order of magnitude of the value added due to the first reason is only $O(\sqrt{t^{\prime}})$, which is asymptotically neglectable in comparison with $t^{\prime}$, it remains to consider only the second reason.
Assume that we can pack in endpoints only details with indices ranging from $t^{\prime}$ to $Ct^{\prime}$, where $C>1$ is some fixed number independent of $t^{\prime}$. But then the total area of these boxes is asymptotically not less than
@@
Really, inequality (4) is equivalent to the fact that the box $B_{n}$ satisfies the inequality $(1+\sigma)/n^{\gamma}>w(B_{n})$, therefore the width of any normal box is less than $(1+\sigma)/n_{0}^{\gamma}$, whence we conclude that $t_{0}>n_{0}^{\gamma}/(1+\sigma)$. Two-sided constraints imply the correlation $t_{0}\sim n_{0}^{\gamma}$.
@@ 5 Proof of the main theorem
$S_{ep,2}(t)=\sum_{B\in{\mathcal{B}}_{ep,2}(t)}w^{2}(B)\frac{h(B)}{w(B)}=O\bigl{(}\frac{1}{t^{2}}\bigr{)}|{\mathcal{B}}_{ep,2}(t)|\times\frac{1}{|{\mathcal{B}}_{ep,2}(t)|}\sum_{B\in{\mathcal{B}}_{ep,2}(t)}\frac{h(B)}{w(B)}.$
@@
In addition, according to formula (7), $S_{com}(t)\sim 1/t$. For $t\leq t_{0}$ the assertion of the theorem is valid by Lemma 8, and there is no need in using assumptions 1 and 2. Let us now consider the general case for all $t\in T_{crit}$. Let us divide the left- and right-hand sides of equality (13) by $S_{com}(t)$. The second addend satisfies inequality (5), the third one does Lemma 4, the fourth one, according to Lemma 10, satisfies the bound $S_{ep,2}(t)/S_{com}(t)=o(1)$. ∎
@@ 6 Conclusion
The graphs given below (see Fig. 8 and Fig. 9) illustrate the dependence of the ratio of LRP on the value of the parameter $n_{0}$ with $\gamma=4/3$, when the Slack-Pack algorithm is being used for packing up to $10^{10}$ details. The horizontal dashed line corresponds to the value of $1-1/\gamma$, while the vertical dash-dotted line does to the argument $n_{0}^{\gamma}\sim t_{0}$. According to experimental results, after attaining the value $t=t_{0}$ the ratio of LRP stabilizes at the level slightly exceeding $1-1/\gamma-\delta$, where $\delta$ is some small value.
@@
We have also performed certain experiments for verifying inequality (4) in Assumption 1. For the value $1+\sigma$ we used the expression $h({\widetilde{B}})^{\gamma}/w({\widetilde{B}})$, where ${\widetilde{B}}$ denotes the last normal box in the first stripe. This choice is motivated by the fact that among all normal boxes of the first kind this particular box yields the smallest value of the ratio $h(B)^{\gamma}/w(B)$. Experiments were performed for sets of details $R_{n}$ with the same values of $n_{0}$ and $\gamma$ as those used in previous experiments. According to obtained results, inequality (4) is fulfilled for all considered normal boxes. See also the Appendix A for theoretical and experimental arguments in favor of Assumption 2.